# Attention-Guided Multi-scale Interaction Network for Face Super-Resolution

### v3

@@ VI Conclusions
This work proposes an attention-guided Multi-scale interaction network for face super-resolution. Specifically, we design an LGFI, which allows accessible communication of global features obtained from self-attention and local features obtained from our designed RDFE. To enhance the variegation of local features in RDFE, we employ a multi-scale depth separable convolutional kernel coupled with attention mechanisms to extract and refine local features. Furthermore, to adaptively fuse features at different scales, we propose an RDFE to utilize the attention mechanism to select a convolutional kernel of the appropriate size to promote feature fusion. Extensive experiments on the synthetics and real test sets show our designed modules significantly improve the communication of features at different scales with modules, allowing our proposed method to outperform existing methods regarding FSR performance, model size, and inference speed.

### v4

@@ VI Conclusions
We propose an attention-guided multiscale interaction network for face super-resolution. The core component, LGFI, facilitates effective interaction between global features from self-attention and local features extracted by the proposed RDFE module. To enrich local representations, RDFE employs multiscale depthwise separable convolutions combined with attention for feature extraction and refinement. Moreover, an adaptive kernel selection mechanism further promotes multiscale feature fusion. Extensive experiments on synthetic and real-world datasets demonstrate that our design substantially enhances cross-scale feature interaction, enabling our method to surpass existing approaches in reconstruction quality, model size, and inference efficiency.